5. On the rod with a length of 70 cm that has a rotating shaft at the end O, the force F = 40N acts, the direction of which forms the angle 60⁰ with the rod. Find the moment of force acting on the rod.


1
Expert's answer
2021-03-26T11:40:53-0400

"\tau=Fl\sin\alpha=40\cdot0.7\cdot\sin60°=24.25\ (N\cdot m)" . Answer